# Nonlinear integral extension of PID control with improved convergence of perturbed second-order dynamic systems

### v1

@@ Abstract
Nonlinear extension of the integral part of PID feedback control is proposed for the perturbed second-order systems. For the matched constant perturbations, the global asymptotic stability is shown, and for Lipschitz perturbations an ultimately bounded output error is guaranteed. The second-order system plants can also be expanded by an additional (parasitic) actuator dynamics with low-pass characteristics. The proposed nonlinear control is proven to outperform its linear (PID) benchmarking counterpart during the settling phase, i.e. at convergence of the residual output error. An experimental case study of the second-order system with an additional actuator dynamics and considerable perturbation is demonstrated to confirm and benchmark the control performance.
I INTRODUCTION
@@
Proportional-integral-derivative (PID) controllers, see e.g. [1, 2], as well as their possible extensions like anti-windup, gain-scheduling, parameters adaptation, are still mostly seen as a working horse in various industries. Despite a large number of available and well perceived tutorials and surveys on PID principles, tuning, extensions, and limitations, e.g. [3, 4, 5] just for mentioning here the few, numerous applications face (here and again) some crucial challenges with use of the PID controllers. These lie often in a trade-off between the performance and robustness, while the metrics associated with a transient and steady-state response appear to be most significant for several applications.
@@
Well known, an integral control action is required for asymptotic set-point regulation under parameter perturbations or (equally usual in practice) presence of the matched stationary disturbances. For simple first- and second-order processes, including possibly also time delays, the tuning of the integral time constant is straightforward and well understood, see e.g. [6]. More challenging becomes an effective integral control action in case of disturbance processes that vary over time, especially in vicinity of the set-point, and when the application requires an increased settling accuracy in shorter time. Not surprising is, hence, that a standard integral control action cannot, for example, compensate accurately for the nonlinear Coulomb friction [7, 8] in the set-point tasks. As one of the possible solutions, the so-called reset integral control (introduced originally in [9]) can be applied, see e.g. [10]. The reset integral behavior in feedback, and thus a switching dynamics, restricts however the applicability to a certain class of the system plants [11]. Worth noting is that despite a relaxation applied to the underlying base systems and, this way, an attempt to address the stability by a less stringent circle-criterion conditions [12], the reset integrator does not entirely satisfy the necessary conditions posed on the feedback nonlinearity, cf. [13, 14], cf. with preliminaries provided below in section II. Another approach of the time-varying integral control gain, which assumes a monotonic decrease of the gain to zero at a sufficiently slow rate, was presented in [15] for linear system plants subject to a globally Lipschitz and non-decreasing actuator nonlinearity. Despite the profound theoretical results, which also form basis for a simple adaptive control strategy, the approach is not directly applicable to compensate for external disturbances and to improve the convergence performance. An universal integral controller for minimum-phase nonlinear systems and the tuning procedures for its key parameters were proposed in [16]. A single-input-single-output (SISO) nonlinear system plant that has a well-defined normal form with asymptotically stable zero dynamics was assumed for the synthesis procedure. Remarkable is that for relative degree-one systems, this universal controller reduces to the classical PI controller followed by saturation, and for relative degree-two systems it reduces to the classical PID controller followed by saturation. Settling performance, in terms of the convergence time and accuracy were, however, not explicitly in focus of the proposed methodology. Later, the problem of robust nonlinear integral control for a class of systems with dynamic uncertainties was addressed in [17], while a constructive controller design procedure was presented in the presence of an unknown equilibrium due to uncertain nonlinearities and dynamic uncertainties. The resulted controller reduces, notwithstanding, to linear control laws if the nominal system is linear, and becomes a traditional PI control if, additionally, the relative degree is one, or the uncertainty falls into the input space, i.e. strict matching condition.
Motivated by the background provided above, and the steady-state behavior of motion control systems, see e.g. [18], the present work introduces a possible nonlinear integral extension of the PID feedback control. To the best of the author’s knowledge, the proposed control, see eq. (8), is new and was not used in the previously published works.
@@
The rest of the paper is structured in the following way. Section II provides the notation in use and the necessary preliminaries of a standard PID feedback control and circle criterion of absolute stability. The main results are placed into section III, starting with absolute stability, followed by the disturbance rejection and extension of the second-order system plant by an additional actuator dynamics. Also some clarifying numerical examples are included. An experimental case study used for evaluation of the proposed control and benchmarking it with the standard PD and PID controllers is provided in section IV. The paper ends with the short conclusions drawn in section V.
II PRELIMINARIES
@@
II-A PID controlled feedback loop
@@
where the output of interest $y(t)$ and its time derivative $\dot{y}(t)$ are available for a feedback control. The control signal $u(t)\in\mathbb{R}$ is continuous on $t\in\mathbb{R}_{+}$ and assumed to stay unsaturated [^1] We notice that a feedback control involving an integral action may be subject to saturation, which is also quite common practice in control engineering. However, an explicit measure against the associated wind-up effects is beyond the scope of the current work. Therefore, neither standard PID control, assumed as a reference benchmarking control system, nor its nonlinear extension proposed handle a saturated operation mode. for $(y,\dot{y})\in\mathbb{R}^{2}$. The matched perturbation function $\sigma(t)$ is assumed to be Lipschitz, i.e. $|\dot{\sigma}|<\Sigma$.
First, we recall briefly a standard PID feedback control
@@
with the design parameters $a,b,c>0$, while focusing below on the most relevant (for us) properties and features. Control parameters are assumed to be appropriately determined.
One can easily show, either in time or frequency domain, that for a constant disturbance, i.e. $\dot{\sigma}=0$, the PID feedback control (2) provides a global asymptotic stabilization of (1). For a Lipschitz disturbance, mapped into frequency domain by $\Sigma\propto\omega$, the PID feedback control (2) provides an ultimately bounded output error $|y(t)|<\varepsilon(\omega)$ for $t\rightarrow\infty$. Regarding this statement, which is rather self-understood in the linear feedback control theory, cf. e.g. [19, 2], we will have a closer look below in section III-B.
@@
one can show by, for example Routh-Hurwitz stability criterion, see e.g. [19], that the closed-loop control system is asymptotically (and exponentially) stable if and only if
@@
II-B Circle criterion
For a linear time invariant (LTI) system in the minimal state-space realization, see e.g. [20] for detail, the nonlinearity $\phi(\cdot,\cdot)$ that is acting in feedback and satisfying
@@ II-B Circle criterion
results in the so called Lur’e problem [21]. The latter is also refereed to as the absolute stability problem and is formulated for a generic class of systems given by
@@
Note that the single-input-single-output (SISO) systems are considered, so that $u,\,z\in\mathbb{R}$, while $A,\,B,\,C,\,D$ are the matrices (correspondingly vectors) of an appropriate dimension and contain the system parameters. Here and for the rest of the section, we will closely follow the notations and developments of circle criterion [22, 23] as these are provided in [14]; for additional explanations see also [24].
For applying the circle criterion of absolute stability to (II-B), the following necessary assumptions are placed.
@@
    $A$, $B$, and $C$ are the minimal realization of a linear subsystem in (II-B), i.e. the pair $(A,B)$ is fully controllable and the pair $(A,C)$ is fully observable.
@@
of the linear subsystem in (II-B), the absolute stability of the closed-loop (II-B) with $\phi(\cdot,t)$ belonging to the sector $[0,k]$, is guaranteed by the circle criterion stated in the below theorem.
Theorem 1
Given a feedback system described by the equations (II-B). Let the assumptions (i)-(iv) be effective. Then, the origin of (II-B) is globally asymptotically stable if
@@
The proof of Theorem 1 can be found in [14, page 241].
It is worth recalling that the circle criterion given by (7) for the sector $[0,k]$ is equivalent to, correspondingly can be transformed into (and also from), the circle criterion formulated for the sector $[k_{1},k_{2}]$ in terms of a disc $\mathcal{D}(k_{1},k_{2})$ in the polar coordinates of $H(j\omega)$-locus, see e.g. [24, 14].
III MAIN RESULTS
The proposed feedback control law with nonlinear extension of the integral part is given by, cf. (2),
@@ III MAIN RESULTS
The linear control parameters $a,b,c>0$ are in the same sense as for (2). The nonlinear parameters $d\geq 0$ and $e<0$ provide a smooth variation of the overall integral gain
@@
that is monotonically increasing with decrease of the magnitude of the control error. Worth emphasizing is that
@@
In the following, we are proving the absolute stability of an unperturbed system (1), (8) by using the circle criterion, cf. with section II-B. Afterwards, the closed-loop system (1), (8) with an exogenous input $\sigma$ is analyzed in terms of the convergence performance of $y(t)$, and that in comparison to the control law (2), i.e. for $d=0$ in (8).
III-A Absolute stability
The unperturbed closed-loop system (1), (8) is transferable into the Lur’e problem (II-B) with $D=0$ and
$A=\left(\begin{array}[]{ccc}0&1&0\\ 0&0&1\\ -c&-b&-a\\ \end{array}\right),\,B=\left(\begin{array}[]{c}0\\ 0\\ 1\\ \end{array}\right),\,C^{\top}=\left(\begin{array}[]{c}0\\ 0\\ 1\\ \end{array}\right).$ (10)
The necessary assumptions i) and ii), cf. with section II-B, are fulfilled since the state-space realization (10) is both fully controllable and observable, and one can guarantee that the system matrix $A$ is Hurwitz under the condition that inequality (4) is satisfied. The feedback nonlinearity
@@ III-A Absolute stability
proves to satisfy the necessary assumption iii), since it belongs to the sector $[0,k]$ with $k=cd>0$ for all $|x_{2}(t)|\in\mathbb{R}_{+}$. Fulfillment of the necessary assumption iv) follows from the fact that (11) is single-valued with respect to the $(z,|x_{2}|)$ pair, i.e. it is memoryless, and continuous in $t$ and Lipschitz in $z$. Two of the last statements are due to the exponential term, and since $x_{2}(t)\in\mathcal{C}^{3}$ for all $u(t)$ and $\sigma(t)$ to be Lipschitz. Now, we are in position to show the absolute stability of the unperturbed control system (1), (8).
Theorem 2
@@
Evaluating the real part of (13) and substituting it, together with $k=cd$, into (7) results in
@@
Comparing it to (12), this completes the proof. ∎
Remark 1
@@
A graphical visualization of the circle stability criterion applied to $H(j\omega)$, with an exemplary stable characteristic polynomial (3) satisfying (12) and $k=cd$, is depicted in Fig. 1 for the sake of explainability.
III-B Disturbance rejection
Consider a standard feedback control loop as shown in Fig. 2. Note that if a reference $r(t)$ is also applied, then both controllers (2) and (8) will incorporate the control error $\epsilon=r-y$ instead of $y$ and, respectively, its derivative and integral. Consequently, they will change to additive inverse, i.e. $-a\dot{y}$ changes to $a\dot{\epsilon}$ and so on. In the following, for the sake of simplicity and due to the focus on disturbance rejection, we assume $r=0$. Same time, $\sigma\neq 0$ is entering the loop. Further, the control (8) is denoted by nl-PID while its linear counterpart (2) by PID, respectively, cf. Fig. 2.
Figure 2: Closed loop system with (nl-)PID control.
While the system plant transfer function is given by
@@ III-B Disturbance rejection
Substituting (14), (15) and evaluating (16), one can recognize that for any $\sigma=\mathrm{const}$, the resulting $y(s)\rightarrow 0$ as $s\rightarrow 0$, i.e. at steady-state. This is irrespectively of the $\Omega$-value for which the characteristic polynomial (3) with $c=\Omega$ is guaranteed to remain stable. Evaluating the magnitude of the frequency response function $S_{\sigma y}(j\omega)$ one obtains
@@
One can show that for an increasing $\Omega$, while satisfying $\Omega<ab$, the peak value of (17) also increases, i.e. $\Omega\nearrow\>\Rightarrow\>\max\,\bigl{|}S_{\sigma y}(j\omega)\bigr{|}\nearrow$. A qualitative behavior of (17) is shown in Fig. 3 for the exemplary assigned $\Omega\in\{1c,2c,3c,4c\}$.
@@
Same time, one can also show that for frequencies lower than the peak frequency $\omega_{\mathrm{peak}}$, an increasing $\Omega$ leads to a decreasing magnitude value, i.e. $\Omega\nearrow\>\Rightarrow\>\bigl{|}S_{\sigma y}(j\omega)\bigr{|}\searrow$ for $\omega\ll\omega_{\mathrm{peak}}$, cf. Fig. 3. Here we recall that higher peak values give rise to higher overshoots during the transients, like for example in case when all frequencies are excited by a stepwise $\sigma$. On the contrary, lower $|S_{\sigma y}|$ values reduce the ultimately bounded output error, correspondingly, improve the convergence performance. Thus, an appropriately varying $\Omega$ value, cf. (9), can improve both the transient and steady-state behavior of the control system. We should recall, however, that the behavior (17), cf. also Fig. 3, discloses only a steady-state response of the feedback control system.
@@
In the following, for the sake of simplicity when developing and analyzing the time domain solution of (18) we assume the real eigenvalues $\lambda_{1},\lambda_{2}>0$ so that the stable transfer function (16) has a double-pole pair at $-\lambda_{1}$ and, respectively, one pole at $-\lambda_{2}$. One can show that such poles configuration corresponds to the control gains equal to
@@
Solving (18) in time domain and putting outside the brackets the common term $\Gamma=(\lambda_{1}-\lambda_{2})^{-2}$ results in
@@
$y(t)$ $=$ $\hat{\Gamma}\biggl{(}\exp\Bigl{(}-\frac{\Omega}{\lambda_{1}^{2}}t\Bigr{)}-\Bigl{(}1-\frac{\Omega-\lambda_{1}^{3}}{\lambda_{1}^{2}}\,t\Bigr{)}\exp\bigl{(}-\lambda_{1}t\bigr{)}\biggr{)},$
$\hat{\Gamma}$ $=$ $\frac{\lambda_{1}^{4}}{(\lambda_{1}^{3}-\Omega)^{2}}.$ (21)
Analyzing (21), one can show that an increasing $\Omega$ (for a fixed $\lambda_{1}$ value) leads to a decreasing peak in the transient response of $y(t)$, cf. Fig. 4 for exemplary $\Omega\in\{0.5c,1c,2c\}$.
@@
The following numerical example visualizes convergence of the control system (1), (8) for various parameter settings. The assigned (most simple) disturbance is a constant value $\sigma=-100$. The set linear control parameters are $a=60$, $b=1100$, $c=3000$, resulting in a poles configuration $\lambda\in\{-3.28,\,-28.36\pm j10.47\}$, i.e. when the nonlinear integral gain is excluded, i.e. $d=0$. Note that for an integral gain equal to $60002c=$ 6000, that corresponds to $d=1$ when $y\rightarrow 0$, the poles configuration results in $\lambda^{*}\in\{-10,\,-20,\ -30\}$. The initial value $y(0)=-1$ is assigned for a simultaneous set-point and disturbance rejection control task.
For a fixed parameter $e=-10$, the output convergence on the logarithmic scale is compared for the set parameter values $d=\{1,2,3\}$ in Fig. 5. The linear PID case, i.e. for $d=0$, is also included for the sake of comparison.
@@
Next, for a fixed parameter $d=2$, the output convergence on the logarithmic scale is compared for the set parameter values $e=\{-10,-100,-1000\}$ in Fig. 6. Also here, the linear PID case of $d=0$ is shown for comparison.
@@
One can recognize that for some lower threshold $\varepsilon=\mathrm{const}$ of the residual error, the convergence of the nl-PID control outperforms that one of the PID, in terms of $|y(t)|<\varepsilon$ for $t>t_{0}$ with $t_{0}$ to be a metric of the settling time. Simultaneously, for higher $\varepsilon$ values, for instance with $t_{0}<1$ sec cf. Figs. 5, 6, it is a combination of $(d,e)$ parameters that allows shaping the output convergence to be superior comparing to the linear case. Note that in addition to $y(t)$ convergence, the transient over- or undershoots and the control effort can appear as essential in a particular application.
III-C Extension by an actuator dynamics

### v3

# MICHAEL RUDERMAN

@@ Abstract
[Summary]Nonlinear extension of the integral part of a standard proportional-integral-derivative (PID) feedback control is proposed for perturbed second-order systems. The approach is model-free and requires solely the Lipschitz boundedness of the unknown matched perturbations. For constant disturbances, the global asymptotic stability is shown based on the circle criterion. For Lipschitz perturbations, an ultimately bounded output error is provided based on the steady-state behavior in frequency domain. Also the transient response to the stepwise disturbances is analyzed for the control tuning. Based on the developed analysis, the design recommendations are formulated as a step by step procedure. It is also discussed how the proposed control is applicable to second-order systems extended by additional (parasitic) actuator dynamics with low-pass characteristics. The proposed nonlinear control is proven to outperform its linear PID counterpart during the settling phase, i.e. at convergence of the residual output error. An experimental case study of the second-order system with an additional actuator dynamics and considerable perturbations is demonstrated to confirm and benchmark the control performance.
keywords:
Nonlinear control, integral feedback, stability analysis, PID control, feedback stabilization
articletype: author’s accepted manuscript by Wiley publisher
1 INTRODUCTION
@@
Proportional-integral-derivative (PID) controllers, see e.g. 3, 4, as well as their possible extensions like anti-windup 13, gain-scheduling, parameters adaptation, 11, and others 22 are still seen as a mostly working horse in various industries. Despite a large number of the available and well perceived tutorials and surveys on the PID principles, tuning, extensions, and limitations, e.g. 1, 12, 10 just for mentioning here the few, numerous applications face here and again some crucial challenges with use of the PID controllers. This is often due to a trade-off between the respondness and robustness, while the performance metrics associated with a transient and steady-state response appear to be most significant for several applications.
@@
Well known, an integral control action is required for asymptotic set-point regulation under parameter perturbations or in presence of the matched stationary disturbances, which are quite common in the control practice. For simple first- and second-order processes, including also the possible time delays, the tuning of the integral time constant is straightforward and well understood, see e.g. 29. More challenging becomes an effective integral control action in case of disturbance processes that vary over time, especially in vicinity to the set-point, and when the application requires an increased settling accuracy in a shorter time. Hence it is not surprising that a standard integral control action cannot compensate accurately, for example, the nonlinear Coulomb friction 7, 24, 25 in the set-point tasks. As one of the possible solutions, the so-called reset integral control, which was introduced originally in 8, can be applied, see e.g. 5. The reset integral behavior in feedback, and thus a switching dynamics, restricts however the applicability to a certain class of the system plants 6. Worth noting is also that despite a relaxation applied to the underlying base systems and, this way, an attempt to address the stability by a less stringent circle-criterion conditions 31, the reset integrator does not entirely satisfy the necessary conditions posed on the feedback nonlinearity, cf. 16, 28, cf. with preliminaries provided below in section 2. Another approach of the time-varying integral control gain, which assumes a monotonic decrease of the gain to zero at a sufficiently slow rate, was presented in 18 for linear system plants subject to a globally Lipschitz and non-decreasing actuator nonlinearity. Despite the profound theoretical results, which also form the basis for a simple adaptive control strategy, the approach is not directly applicable to compensate for external disturbances and to improve the convergence performance. An universal integral controller for minimum-phase nonlinear systems and the tuning procedures for its key parameters were proposed in 15. A single-input-single-output (SISO) nonlinear system plant that has a well-defined normal form with asymptotically stable zero dynamics was assumed for the synthesis procedure. Remarkable is that for relative degree-one systems, this universal controller reduces to the classical PI controller followed by the saturation, and for relative degree-two systems it reduces to the classical PID controller followed by the saturation. Settling performance, in terms of the convergence time and accuracy were, however, not explicitly in focus of the proposed methodology. Later, the problem of robust nonlinear integral control for a class of systems with dynamic uncertainties was addressed in 14, while a constructive controller design procedure was presented in the presence of an unknown equilibrium due to uncertain nonlinearities and dynamic uncertainties. The resulted controller reduces, notwithstanding, to linear control laws if the nominal system is linear, and becomes a traditional PI control if, additionally, the relative degree is one, or the uncertainty falls into the input space, i.e. it fulfills the strict matching condition.
Motivated by the background provided above and the perturbed steady-state behavior of motion control systems, see e.g. 23, the present work introduces a possible nonlinear integral extension of the PID feedback control. To the best of the author’s knowledge, the proposed control, see eq. (8), is new and was not reported in the previously published works. An innovative contribution of the work lies in extending a standard linear PID feedback controller by a nonlinear integral part which allows for faster output convergence in presence of the unknown matched Lipschitz perturbations. Such extension is model-free and relies solely on the Lipschitz continuity of the disturbing perturbations. A further contribution is in the simple analysis of stability and output error dynamics as well as in the illustrative experimental case study that supports the proposed control approach.
@@
The rest of the paper is structured in the following way. Section 2 provides the notation in use and the necessary preliminaries of a standard PID feedback control and the circle criterion of the absolute stability. The main results are placed into section 3, starting with absolute stability, followed by the disturbance rejection and extension of the second-order system plant by an additional actuator dynamics. Also some clarifying numerical examples are included, and a step-by-step controller design procedure is recommended. An experimental case study used for evaluation of the proposed control and benchmarking it against the standard PD and PID controllers is provided in section 4. The paper ends with the short conclusions drawn in section 5.
2 PRELIMINARIES
@@
2.1 PID controlled feedback loop
@@
where the output of interest $y(t)$ and its time derivative $\dot{y}(t)$ are available for a feedback control. The control signal $u(t)\in\mathbb{R}$ is continuous on $t\in\mathbb{R}_{+}$ and assumed to stay unsaturated [^1] We notice that a feedback control involving an integral action may be subject to saturation, that is quite common in practice of the control engineering. However, an explicit measure against the associated wind-up effects is beyond the scope of the current work. Therefore, neither standard PID control, assumed as a reference benchmarking control system, nor the proposed nonlinear extension handle a saturated operation mode. for $(y,\dot{y})\in\mathbb{R}^{2}$. The matched perturbation function $\sigma(t)$ is assumed to be Lipschitz, i.e. $|\dot{\sigma}|<\Sigma$ for some positive known $\Sigma$.
First, we recall briefly a standard PID feedback control, cf. e.g. 3, 4, which in case of zero reference has the form
@@
with the design parameters $a,b,c>0$. Below we are focusing on the most relevant (for us) properties and features of the PID control, while the control parameters are assumed to be appropriately determined.
One can easily show, either in time or frequency domain, that for a constant disturbance, i.e. $\dot{\sigma}=0$, the PID feedback control (2) provides a global asymptotic stabilization of (1). For a Lipschitz disturbance, mapped into frequency domain by $\Sigma\propto\omega$, the PID feedback control (2) provides an ultimately bounded output error $|y(t)|<\varepsilon(\omega)$ for $t\rightarrow\infty$. Regarding this statement, which is rather self-understood in the linear feedback control theory, cf. e.g. 9, 4, we will have a closer look below in section 3.2.
@@
one can show by, for example Routh-Hurwitz stability criterion, e.g. 9, that the closed-loop control system is asymptotically (and exponentially) stable if and only if
@@
2.2 Circle criterion
For a linear time invariant (LTI) system in the minimal state-space realization, see e.g. 2 for detail, a nonlinearity $\phi(\cdot,\cdot)$ that is acting in feedback and satisfying
@@ 2.2 Circle criterion
results in the so called Lur’e problem 19. The latter is also refereed to as the absolute stability problem and is formulated for a generic class of the systems given by
@@
Note that the SISO systems are considered, so that $u,\,z\in\mathbb{R}$, while $A,\,B,\,C,\,D$ are the matrices (correspondingly vectors) of an appropriate dimension, and they contain the system parameters. Here and for the rest of the section, we will closely follow the notations and developments of the circle criterion 21, 27 as it is provided in 28; for additional explanations see also e.g. 30.
For applying the circle criterion of absolute stability to (2.2), the following necessary assumptions are placed.
@@
    $A$, $B$, and $C$ are the minimal realization of a linear subsystem in (2.2), i.e. the pair $(A,B)$ is fully controllable and the pair $(A,C)$ is fully observable.
@@
of the linear subsystem in (2.2), the absolute stability of the closed-loop (2.2) with the nonlinearity $\phi(\cdot,t)$ belonging to the sector $[0,k]$, is guaranteed by the circle criterion stated in the following theorem.
Theorem 1.
Given a feedback system described by the equations (2.2). Let the assumptions (i)-(iv) be effective. Then, the origin of (2.2) is globally asymptotically stable if
@@
The proof of Theorem 1 can be found in 28 page 241.
It is worth recalling that the circle criterion given by (7) for the sector $[0,k]$ is equivalent to, correspondingly can be transformed into (and also from), the circle criterion formulated for the sector $[k_{1},k_{2}]$ in terms of a disc $\mathcal{D}(k_{1},k_{2})$, which is drawn in the polar coordinates of the $H(j\omega)$-locus, see e.g. 30, 28.
3 MAIN RESULTS
The proposed feedback control law with a nonlinear extension of the integral part is given by
@@ 3 MAIN RESULTS
cf. (2). The linear control parameters $a,b,c>0$ are in the same sense as for (2). The nonlinear parameters $d\geq 0$ and $e<0$ provide a smooth variation of the overall integral gain
@@
that is monotonically increasing with decrease of the magnitude of the control error. It is worth emphasizing that:
@@
In the following, we are first proving the absolute stability of an unperturbed system (1), (8) by using the circle criterion, cf. with section 2.2. Afterwards, the closed-loop system (1), (8) with an exogenous input $\sigma$ is analyzed in terms of the convergence performance of $y(t)$, and that in comparison to the control law (2), i.e. for $d=0$ in (8). An illustrative numerical example discloses the output convergence properties in presence of a constant disturbance when varying $d$-amplitude. Here we recall that the integral control part (of a standard PID control) is expected to compensate for constant unknown disturbances at steady-state. Then, a system extension by the additional (parasitic) actuator dynamics is also addressed. Finally, we draw several recommendation lines for design (in terms of the parameters tuning) of the proposed nonlinear extended PID controller.
3.1 Absolute stability
The unperturbed closed-loop system (1), (8) is transferable into the Lur’e problem (2.2) with $D=0$ as
$A=\left(\begin{array}[]{ccc}0&1&0\\ 0&0&1\\ -c&-b&-a\\ \end{array}\right),\,B=\left(\begin{array}[]{c}0\\ 0\\ 1\\ \end{array}\right),\,C^{\top}=\left(\begin{array}[]{c}1\\ 0\\ 0\\ \end{array}\right),$ (10)
[…]
@@ 3.1 Absolute stability
proves to satisfy the necessary assumption iii), since it belongs to the sector $[0,k]$ with $k=cd>0$ for all $|x_{2}(t)|\in\mathbb{R}_{+}$. Fulfillment of the necessary assumption iv) follows from the fact that (11) is single-valued with respect to a $(z,|x_{2}|)$ pair, i.e. it is memoryless, and continuous in $t$ and Lipschitz in $z$. Two of the last statements are true due to the exponential term, and since $x_{2}(t)\in\mathcal{C}^{3}$ for all $u(t)$ and $\sigma(t)$ to be Lipschitz. Now, we are in the position to show the absolute stability of the unperturbed control system (1), (8).
Theorem 2.
@@
Evaluating the real part of (13) and substituting it together with $k=cd$ into (7) results in
@@
By comparing it with (12), the proof is completed. ∎
Remark 1.
@@
A graphical visualization of the circle stability criterion applied to $H(j\omega)$, with an exemplary stable characteristic polynomial (3) which is satisfying (12) and $k=cd$, is depicted in Figure 1 for the sake of a qualitative explainability.
3.2 Disturbance rejection
Consider a standard feedback control loop as shown in Figure 2. Note that if a reference $r(t)\neq 0$ is also applied, then both controllers (2) and (8) will incorporate the control error $\epsilon=r-y$ instead of $y$ and, respectively, its derivative and integral. Recall that this is a rather standard step of a coordinates transformation, since a stabilization problem (respectively disturbance rejection problem) with $r=0$ was considered before. Consequently, the control terms will change to their additive inverse, i.e. $-a\dot{y}$ will change to $a\dot{\epsilon}$ and so on. In the following, for the sake of simplicity and due to the focus on disturbance rejection, we further assume $r=0$. At the same time, an $\sigma\neq 0$ value is entering the control loop. Further, the control (8) will be denoted by nl-PID while its linear counterpart (2) by PID, respectively, cf. Figure 2.
Figure 2: Closed loop system with (nl-)PID controllers.
While the transfer function of the system plant is given by
@@ 3.2 Disturbance rejection
Substituting (14), (15) and evaluating (16), one can recognize that for any $\sigma=\mathrm{const}$, the resulting $y(s)\rightarrow 0$ as $s\rightarrow 0$, i.e. for the steady-state. This is irrespectively of the $\Omega$-value for which the characteristic polynomial (3) with $c=\Omega$ is maintained to remain stable. Evaluating the magnitude of the frequency response function $S_{\sigma y}(j\omega)$ one obtains
@@
One can show that for an increasing $\Omega$, while satisfying $\Omega<ab$, the peak value of (17) also increases, i.e. $\Omega\nearrow\>\Rightarrow\>\max\,\bigl{|}S_{\sigma y}(j\omega)\bigr{|}\nearrow$. A qualitative behavior of (17) is shown in Figure 3 for the exemplary assigned $\Omega\in\{1c,2c,3c,4c\}$.
@@
At the same time, one can also show that for frequencies lower than the peak frequency $\omega_{\mathrm{peak}}$, an increasing $\Omega$ leads to a decreasing magnitude value, i.e. $\Omega\nearrow\>\Rightarrow\>\bigl{|}S_{\sigma y}(j\omega)\bigr{|}\searrow$ for $\omega\ll\omega_{\mathrm{peak}}$, cf. Figure 3. Here we recall that higher peak values give rise to higher overshoots during the transients, like for example in case when all frequencies will be excited in case of a stepwise $\sigma$ disturbance. On the contrary, lower $|S_{\sigma y}|$ values reduce the ultimately bounded output error, correspondingly, improve the convergence performance. Thus, an appropriately dimensioned $\Omega$-value, cf. (9), can improve both the transient and steady-state behavior of the control system. We should consciously recall, however, that the behavior (17), cf. also Figure 3, discloses only a steady-state response of the feedback control system.
@@
In the following, for the sake of simplicity of developing and analyzing the time domain solution of (18), we assume two real eigenvalues $\lambda_{1},\lambda_{2}<0$ so that the stable transfer function (16) has a double-pole pair at $-\lambda_{1}$ and, respectively, one pole at $-\lambda_{2}$. One can show that such poles configuration corresponds to the control gains equal to
@@
Solving (18) in time domain and putting the common term $\Gamma=(\lambda_{1}-\lambda_{2})^{-2}$ outside the brackets results in
@@
$y(t)$ $=$ $\hat{\Gamma}\biggl{(}\exp\Bigl{(}-\frac{\Omega}{\lambda_{1}^{2}}t\Bigr{)}-\Bigl{(}1-\frac{\Omega-\lambda_{1}^{3}}{\lambda_{1}^{2}}\,t\Bigr{)}\exp\bigl{(}-\lambda_{1}t\bigr{)}\biggr{)},$ (21)
$\hat{\Gamma}$ $=$ $\frac{\lambda_{1}^{4}}{(\lambda_{1}^{3}-\Omega)^{2}}.$
Analyzing (21), one can show that an increasing $\Omega$ (this for a fixed $\lambda_{1}$ value) leads to a decreasing peak in the transient response of $y(t)$, cf. Figure 4 for the exemplary taken $\Omega\in\{0.5c,1c,2c\}$.
@@
The following numerical example visualizes the convergence of the control system (1), (8) for various parameter settings. The assigned (most simple) disturbance is a constant value $\sigma=-100$. The set linear control parameters are $a=60$, $b=1100$, $c=3000$. This is resulting in a poles configuration $\lambda\in\{-3.28,\,-28.36\pm j10.47\}$ when the nonlinear integral gain is excluded, i.e. $d=0$. Note that for an integral gain equal to $60002c=$ 6000, that corresponds to $d=1$ when $y\rightarrow 0$, the poles configuration results in $\lambda^{*}\in\{-10,\,-20,\ -30\}$, thus without transient oscillations in the controlled response. The initial value $y(0)=-1$ is assigned for a simultaneous set-point and disturbance rejection control task.
For the assigned parameter $e=-10$, the output convergence on the logarithmic scale are compared for the set parameter values $d=\{1,2,3\}$ in Figure 5. The linear PID case, i.e. for $d=0$, is also included for the sake of comparison.
@@
Next, for the fixed parameter $d=2$, the output convergence on the logarithmic scale is compared for the set parameter values $e=\{-10,-100,-1000\}$ in Figure 6. Also here, the linear PID case of $d=0$ is also shown for the sake of comparison.
@@
One can recognize that for some lower threshold $\varepsilon=\mathrm{const}$ of the residual output error, the convergence of the nl-PID control outperforms that one of the PID, in terms of $|y(t)|<\varepsilon$ for $t>t_{0}$ with $t_{0}$ to be a metric of the settling time. Simultaneously, for higher $\varepsilon$ values, for instance with $t_{0}<1$ sec cf. Figures 5, 6, it is a combination of the $(d,e)$ parameters that allows shaping the output convergence to be superior comparing to the linear control case. Note that in addition to the $y(t)$-convergence, the transient over- or undershoots and the control effort can appear as essential performance metrics for a particular application at hand. A further explicit analysis of this, is however out of scope of the recent communication.
3.3 Controller design
For use of the nl-PID control developed above, the controller design recommendations are summarized as follows.
  1. 1.
    Design a standard PID feedback controller (2) with respect to the application-specific reference following and disturbance rejecting performance criteria, see e.g. 9 4. For the simple dynamic processes, the analytic tuning rules can be found e.g. in 29. For an optimal disturbance rejection in the stable second-order plants by using a disturbance sensitivity function criteria, the procedure proposed in 26 can for example be used. For a detailed treatment of the PID control design, the seminal textbook e.g. 3 is recommended to be consulted.
  2. 2.
    Using the Lipschitz continuity of the perturbations, i.e. the $\Sigma$ upper bound, and the corresponding frequency characteristics of the disturbance sensitivity function (17), determine the application suitable $\Omega$-value, cf. Figure 3.
  3. 3.
    Adjust the $\Omega$-value, if necessary, in order to fulfill the transient response criteria according to (20).
  4. 4.
    For the determined $\Omega$-value, assign the $d$ parameter so that to comply with the property (i), given by (9), and with the stability criteria (12) of Theorem 2.
  5. 5.
    Adjust the $e$ parameter so that to achieve the desired convergence behavior for a given application.
  6. 6.
    Repeat the steps 2.–5. upon the numerical and experimental control evaluation if necessary.
3.4 Extension by an actuator dynamics